a) The rear window in a car is approximately a rectangle, 1.22 m wide and 0.230 m high. The inside rear-view mirror is 0.480 m from the driver's eyes, and 1.41 m from the rear window. What is the minimum height of the rear-view mirror if the driver is to be able to see the entire height of the rear window in the mirror without moving her head?

b) Calculate the minimum width of the rear-view mirror if the driver is to be able to see the entire width of the rear window in the mirror without moving her head.
Answer #1

considering the geometry of similar triangles

h/d = h' / ( L+d)

h = 0.23* 0.48 / ( 1.41 + 0.48)

h = 0.0584 m

======

similarly for width

w= 1.22* 0.48 / ( 0.48 / 1.41)

w = 0.30984 m

======